OpenStudy (anonymous):

Antonio’s weekly allowance is given by the equation A 0.5c 10, where c is the number of chores he does. If he received $16 in allowance one week, how many chores did he do? Select one: a. 15 b. 14 c. 10 d. 12

OpenStudy (anonymous):

The equation is \[A =.5c +10\] correct?

OpenStudy (anonymous):

thats what i thou

OpenStudy (anonymous):

OK. Let A = 16 If so, then subtract 10 from each side.\[16 - 10 = .5c + 10 - 10\] becomes\[6 = .5c\] divide both sides by .5 (or multiply each side by 2, it's the same thing). \[6(.5) = \frac{ .5c }{ .5 }\] So: \[12 = c\]

OpenStudy (anonymous):

okay same what i was thinking